# Synergizing Large Language Models and Task-specific Models for Time Series Anomaly Detection

### v3

@@ 4.2 Prediction Accuracy
Precision, recall, and F1 score are calculated for each subset, and the average results are presented in Tab. 2, where the best performance is highlighted in bolded, and if CoLLaTe achieves the best performance we underline the best performance among baselines. In the table, “prec” and “rec” denote precision and recall, while “OmniAnom” and “AnomalyTr” represent “OmniAnomaly” and “AnomalyTransformer”, respectively. As sigLLM is specifically designed for univariate time series, it cannot be applied to the Mustang, Flight 1 and Flight 2 datasets, and therefore, its performance on these datasets is not included. As shown in Tab. 2, CoLLaTe achieves the highest F1 scores across all four datasets. In this table, GPT-4 represents the LLM method employing the set-up-pitch prompt, which is the LLM component of CoLLaTe. By comparing the performance of GPT-4 with other LLM-based anomaly detection methods, such as sigLLM and LLMAD, it becomes evident that the set-up-pitch prompt significantly enhances the anomaly detection accuracy of the LLM. In Fig. 4(a), we plot the precision and recall corresponding to the best F1 scores as coordinates to illustrate the performance distribution of CoLLaTe, GPT-4, and CoLLaTe. The figure reveals that both the LLM and the TSADM (CoLLaTe) tend to suffer from either low precision or low recall. In contrast, CoLLaTe demonstrates strong performance on both metrics. This observation confirms that CoLLaTe enables effective collaboration between the TSADM and the LLM, resulting in superior overall performance.
@@ 5 Related work
Anomaly detection methods generally fall into two categories: classical anomaly detection methods and neural network-based methods. Classical approaches primarily utilize statistical models (Eskin, 2000; Wang et al., 2016), linear models (Ma & Perkins, 2003; Amer et al., 2013; Hu et al., 2003), and proximity analysis (Li et al., 2021; Izakian & Pedrycz, 2014) to identify anomalies. Nevertheless, these methods can not effectively leverage the expert knowledge.
Neural network-based methods, on the other hand, can be categorized into prediction-based methods (Malhotra et al., 2015; Hundman et al., 2018; Zong et al., 2018; Chen et al., 2022a), reconstruction-based methods (Chen et al., 2022b; You et al., 2022; Jiang et al., 2022; Shen et al., 2021; Tian et al., 2019), and LLM-based methods (Liu et al., 2024b; Russell-Gilbert et al., 2024; Liu et al., 2024a). Prediction-based methods identify anomalies by comparing predicted values to ground truth, but they are vulnerable to historical inference errors (Wang et al., 2024). Reconstruction-based methods address this limitation. However, they cannot leverage expert knowledge in specific application scenarios without model adaptation. These methods identify anomalies by learning normal patterns from training data and flagging data samples that deviate from these patterns. Yet, in scenarios where it is impossible to observe or collect all normal patterns (e.g., airplane monitoring), reconstruction-based methods are ineffective. In contrast, LLM-based methods introduce expert knowledge effectively through techniques like retrieval-augmented generation (RAG) (Lewis et al., 2020) and chain of thought (COT) (Wei et al., 2022) without requiring model modifications. However, as natural language-based models, LLMs are insensitive to value fluctuations and normal pattern extraction. None of these existing methods fully leverage the complementary strengths of LLM-based approaches and TSADM. To address this gap, we propose CoLLaTe, a novel framework designed to enhance anomaly detection performance by combining the advantages of both approaches.
@@ 6 Conclusion
7 Impact Statements
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.
@@ A.1 Proof of Theorem 1
$\begin{split}\min.&-\frac{1}{n}\sum_{i=1}^{n}\log(f(\mathcal{M}(s_{i})))\\ s.t.&\qquad\frac{1}{n}\sum_{k=1}^{n}\mathcal{M}(s_{k})=\hat{\mu}\\ &\frac{1}{n-1}\sum_{k=1}^{n}[\mathcal{M}(s_{k})-\mu_{M}]^{2}=\hat{\sigma}^{2}\end{split}$ (12)
Then, we obtain Eq. 12. After that, by using Lagrange Multiplier method, we can transform Eq. 12 to Eq. 13, where $\hat{\lambda}_{1}>0,\hat{\lambda}_{2}>0$.
$\begin{split}\min.&-\frac{1}{n}\sum_{i=1}^{n}\log(f(\mathcal{M}(s_{i})))+\hat{\lambda}_{1}(\frac{1}{n}\sum_{i=1}^{n}\mathcal{M}(s_{i})-\hat{\mu})^{2}\\ &+\hat{\lambda}_{2}(\frac{1}{n-1}\sum_{i=1}^{n}(\mathcal{M}(s_{i})-\mu_{M})^{2}-\hat{\sigma}^{2})^{2}\end{split}$ (13)
@@ A.2 Proof of Theorem 2
When $\mathcal{L}(a,b)=(a-b)^{2}$, we can transform Eq. 8 to Eq. 14, where $\lambda_{1}=\frac{D_{intra}}{D_{intra}+D_{inter}}$ and $\lambda_{2}=\frac{D_{inter}}{D_{intra}+D_{inter}}$. According to KKT condition (Boyd & Vandenberghe, 2004), the optimal solution of $\mathcal{L}_{\beta}$ should satisfy Eq. 15. As shown in Appendix A.8, the gradient of architecture of conditional network can not be equal to 0. Thus, $02\lambda_{1}(y+\epsilon_{s}-\hat{S}^{*}(\mathcal{P})+2\lambda_{2}(y+\epsilon_{S}-\hat{S}^{*}(\mathcal{P})))=$ 0. Then, $\hat{S}^{*}(\mathcal{P})=y+\lambda_{1}\epsilon_{s}+\lambda_{2}\epsilon_{S}$. Thus, $\mathbb{E}[(\hat{S}^{*}(\mathcal{P})-y)^{2}]=\mathbb{E}[(\lambda_{1}\epsilon_{s}+\lambda_{2}\epsilon_{S})^{2}]$. Since $f(x)=x^{2}$ is a convex function, according to Jensen’s Inequality, $\mathbb{E}[(\lambda_{1}\epsilon_{s}+\lambda_{2}\epsilon_{S})^{2}]\geq[\mathbb{E}(\lambda_{1}\epsilon_{s}+\lambda_{2}\epsilon_{S})]^{2}$. Since $\epsilon_{1}\sim\mathcal{D_{s}}(\mu_{s},\sigma_{s}),\epsilon_{2}\sim\mathcal{D}_{S}(\mu_{S},\sigma_{S})$, $[\mathbb{E}(\lambda_{1}\epsilon_{s}+\lambda_{2}\epsilon_{S})]^{2}=(\lambda_{1}\mu_{1}+\lambda_{2}\mu_{2})^{2}$. Thus, $\mathbb{E}[(\hat{S}^{*}(\mathcal{P})-y)^{2}]\geq(\lambda_{1}\mu_{1}+\lambda_{2}\mu_{2})^{2}$
@@ A.3 Proof of Lemma 1
When using stochastic gradient descent algorithm, the gradient of Eq. 8 and Eq. 9 is shown in Eq. 16 and Eq. 17 respectively, where $\mathcal{B}$ is the set of sample in batch and $|\mathcal{B}|$ is the size of set $\{(r,t)|r\in\mathcal{B},t\in\mathcal{B}\}$. Let $\mathbf{g}_{i,j}^{*}$ denote the item in Eq. 17, whose index is $(i,j)$. Let $\mathbf{g}_{i,j}$ denote the item in Eq. 16, whose index is $(i,j)$. $\mathbf{g}_{i,j}$ can be transformed to $\mathbf{g}_{i,j}^{*}+[\lambda_{1}(\epsilon_{s,i}-\epsilon_{s,j})+\lambda_{2}(\epsilon_{S,i}-\epsilon_{S,j})](\hat{S}_{i}-\hat{S}_{j})\frac{\partial(\hat{S}_{i}-\hat{S}_{j})}{\partial\mathcal{P}}$. Thus, the gradient of Eq. 8 can be deemed as the gradient of Eq. 9 add some random noise scaled by $(\hat{S}_{i}-\hat{S}_{j})\frac{\partial(\hat{S}_{i}-\hat{S}_{j})}{\partial\mathcal{P}}$ for each sample. Let $\hat{S}^{{\dagger}}$ denote the optimal solution of Eq. 9. Thus, $\mathcal{L}_{\beta}^{*}(\hat{S})\geq\mathcal{L}_{\beta}^{*}(\hat{S}^{{\dagger}}),\forall\hat{S}$, and it satisfies Assumption 5.1 in (Bu et al., 2024).
@@
$\mathbf{g}=-\frac{1}{|\mathcal{B}|}\sum_{(i,j)\in\{(r,t)|r\in\mathcal{B},t\in\mathcal{B}\}}\lambda_{1}(\epsilon_{s,i}-\epsilon_{s,j}+y_{i}-y_{j})(\hat{S}_{i}-\hat{S}_{j})\frac{\partial(\hat{S}_{i}-\hat{S}_{j})}{\partial\mathcal{P}}+\lambda_{2}(\epsilon_{S,i}-\epsilon_{S,j}+y_{i}-y_{j})(\hat{S}_{i}-\hat{S}_{j})\frac{\partial(\hat{S}_{i}-\hat{S}_{j})}{\partial\mathcal{P}}$ (16)
$\mathbf{g}^{*}=-\frac{1}{|\mathcal{B}|}\sum_{(i,j)\in\{(r,t)|r\in\mathcal{B},t\in\mathcal{B}\}}\lambda_{1}(y_{i}-y_{j})(\hat{S}_{i}-\hat{S}_{j})\frac{\partial(\hat{S}_{i}-\hat{S}_{j})}{\partial\mathcal{P}}+\lambda_{2}(y_{i}-y_{j})(\hat{S}_{i}-\hat{S}_{j})\frac{\partial(\hat{S}_{i}-\hat{S}_{j})}{\partial\mathcal{P}}$ (17)
@@
Since $y\in[0,1],\hat{S}\in[0,1]$, $|y_{i}-y_{j}|\leq 1,\forall(i,j)$ and $|\hat{S}_{i}-\hat{S}_{j}|\leq 1,\forall(i,j)$. Thus, $|\mathbf{g}_{i,j}^{*}|=|(y_{i}-y_{j})(\hat{S}_{i}-\hat{S}_{j})\frac{\partial(\hat{S}_{i}-\hat{S}_{j})}{\partial\mathcal{P}}|\leq|\frac{\partial(\hat{S}_{i}-\hat{S}_{j})}{\partial\mathcal{P}}|$. Furthermore, according to Appendix. A.8, conditional network is a multi-layer perceptron with sigmoid as its activation function. Let $\mathbf{\sigma}(x)$ denote sigmoid function. Since $\frac{\partial\mathbf{\sigma}(x)}{x}=\mathbf{\sigma}(x)(1-\mathbf{\sigma}(x))$ and $10<\mathbf{\sigma}(x)<$ 1, $|\frac{\partial(\hat{S})}{\partial\mathcal{P}}|<1$. Thus, $|\frac{\partial(\hat{S}_{i}-\hat{S}_{j})}{\partial\mathcal{P}}|\leq 1$. Thus, $|\mathbf{g}_{i,j}^{*}|\leq 1$. Thus, $|\mathbf{g}^{*}|\leq\frac{1}{|\mathcal{B}|}\sum_{i,j}|\mathbf{g}_{i,j}^{*}|=1$ Since $\mathcal{L}_{\beta}^{*}$ is $L$-Lipschitz continuous, $|\mathcal{L}_{\beta}^{*}(\mathcal{P}_{1})-(\mathcal{L}_{\beta}^{*}(\mathcal{P}_{2})+{\mathbf{g}^{*}}^{T}(\mathcal{P}_{1}-\mathcal{P}_{2}))|=|[\mathcal{L}_{\beta}^{*}(\mathcal{P}_{1})-\mathcal{L}_{\beta}^{*}(\mathcal{P}_{2})]+{\mathbf{g}^{*}}^{T}(\mathcal{P}_{1}-\mathcal{P}_{2})|\leq L|\mathcal{P}_{1}-\mathcal{P}_{2}|+|\mathbf{g}_{i,j}^{*}|\cdot|\mathcal{P}_{1}-\mathcal{P}_{2}|\leq(L+1)\cdot|\mathcal{P}_{1}-\mathcal{P}_{2}|$. Since $L\geq 0$, $L+1>0$, it satisfies Assumption 5.2 in (Bu et al., 2024).
@@
Since $\mathbb{E}(\mathbf{g}_{i,j}^{*}-\mathbf{g}_{i,j})=\mathbb{E}[(\lambda_{1}(\epsilon_{s,i}-\epsilon_{s,j})+\lambda_{2}(\epsilon_{S,i}-\epsilon_{S,j}))(\hat{S}_{i}-\hat{S}_{j})\frac{\partial(\hat{S}_{i}-\hat{S}_{j})}{\partial\mathcal{P}}]=\mathbb{E}[(\lambda_{1}(\epsilon_{s,i}-\epsilon_{s,j})+\lambda_{2}(\epsilon_{S,i}-\epsilon_{S,j}))](\hat{S}_{i}-\hat{S}_{j})\frac{\partial(\hat{S}_{i}-\hat{S}_{j})}{\partial\mathcal{P}}$. Since $\mathbb{E}(\epsilon_{s,i})=\mathbb{E}(\epsilon_{s,j}),\mathbb{E}(\epsilon_{S,i})=\mathbb{E}(\epsilon_{S,j})$, $\mathbb{E}[\lambda_{1}(\epsilon_{s,i}-\epsilon_{s,j})+\lambda_{2}(\epsilon_{S,i}-\epsilon_{S,j})]=0$. Thus, $\mathbb{E}(\mathbf{g}_{i,j}^{*}-\mathbf{g}_{i,j})=0$. $\mathbb{E}|\mathbf{g}_{i,j}^{*}-\mathbf{g}_{i,j}|^{2}=(\hat{S}_{i}-\hat{S}_{j})^{2}(\frac{\partial(\hat{S}_{i}-\hat{S}_{j})}{\partial\mathcal{P}})^{2}\mathbb{E}[(\lambda_{1}(\epsilon_{s,i}-\epsilon_{s,j})+\lambda_{2}(\epsilon_{S,i}-\epsilon_{S,j}))^{2}]$. As proven above, $|\hat{S}_{i}-\hat{S}_{j}|\leq 1$, thus $(\hat{S}_{i}-\hat{S}_{j})^{2}\leq 1$. Since $\frac{\partial(\hat{S}_{i}-\hat{S}_{j})}{\partial\mathcal{P}}\leq 1$, $(\frac{\partial(\hat{S}_{i}-\hat{S}_{j})}{\partial\mathcal{P}})^{2}\leq 1$. Thus, $(\hat{S}_{i}-\hat{S}_{j})^{2}(\frac{\partial(\hat{S}_{i}-\hat{S}_{j})}{\partial\mathcal{P}})^{2}\mathbb{E}[(\lambda_{1}(\epsilon_{s,i}-\epsilon_{s,j})+\lambda_{2}(\epsilon_{S,i}-\epsilon_{S,j}))^{2}]\leq\mathbb{E}[(\lambda_{1}(\epsilon_{s,i}-\epsilon_{s,j})+\lambda_{2}(\epsilon_{S,i}-\epsilon_{S,j}))^{2}]$. Since $\epsilon_{s,i}\sim\mathcal{D}(\mu_{s},\sigma_{s}),\epsilon_{s,j}\sim\mathcal{D}(\mu_{s},\sigma_{s})$ independently, $\mathbf{E}[(\epsilon_{s,i}-\epsilon{s,j})^{2}]=2\sigma_{s}^{2}$. Similarly, $\mathbf{E}[(\epsilon_{S,i}-\epsilon_{S,j})^{2}]=2\sigma_{S}^{2}$. After simplifying its formation, we obtain $\mathbb{E}[(\lambda_{1}(\epsilon_{s,i}-\epsilon_{s,j})+\lambda_{2}(\epsilon_{S,i}-\epsilon_{S,j}))^{2}]=2\lambda_{1}^{2}\sigma_{s}^{2}+2\lambda_{2}^{2}\sigma_{S}^{2}$. Thus, $\mathbb{E}|\mathbf{g}_{i,j}^{*}-\mathbf{g}_{i,j}|^{2}\leq 2\lambda_{1}^{2}\sigma_{s}^{2}+2\lambda_{2}^{2}\sigma_{S}^{2}$. Then, it satisfies Assumption 5.3 in (Bu et al., 2024).

### v4

@@ 3.2 Prediction Accuracy
Precision, recall, and F1 score are calculated for each subset, and the average results are presented in Tab. 2, where the best performance is highlighted in bolded, and if CoLLaTe achieves the best performance we underline the best performance among baselines. In the table, “prec" and “rec" denote precision and recall, while “OmniAnom" and “AnomalyTr" represent “OmniAnomaly" and “AnomalyTransformer", respectively. As sigLLM is specifically designed for univariate time series, it cannot be applied to the Mustang, Flight 1 and Flight 2 datasets, and therefore, its performance on these datasets is not included. As shown in Tab. 2, CoLLaTe achieves the highest F1 scores across all four datasets. In this table, GPT-4 represents the LLM method employing the set-up-pitch prompt, which is the LLM component of CoLLaTe. By comparing the performance of GPT-4 with other LLM-based anomaly detection methods, such as sigLLM and LLMAD, it becomes evident that the set-up-pitch prompt significantly enhances the anomaly detection accuracy of the LLM. In Fig. 3(a), we plot the precision and recall corresponding to the best F1 scores as coordinates to illustrate the performance distribution of CoLLaTe, GPT-4, and CoLLaTe. The figure reveals that both the LLM and the TSADM (CoLLaTe) tend to suffer from either low precision or low recall. In contrast, CoLLaTe demonstrates strong performance on both metrics. This observation confirms that CoLLaTe enables effective collaboration between the TSADM and the LLM, resulting in superior overall performance.
@@ 4 Related work
Anomaly detection methods generally fall into two categories: classical anomaly detection methods and neural network-based methods. Classical approaches primarily utilize statistical models Eskin, (2000); Wang et al., (2016), linear models Ma and Perkins, (2003); Amer et al., (2013); Hu et al., (2003), and proximity analysis Li et al., (2021); Izakian and Pedrycz, (2014) to identify anomalies. Nevertheless, these methods can not effectively leverage the expert knowledge.
Neural network-based methods, on the other hand, can be categorized into prediction-based methods Malhotra et al., (2015); Hundman et al., (2018); Zong et al., (2018); Chen et al., 2022a , reconstruction-based methods Chen et al., 2022b ; You et al., (2022); Jiang et al., (2022); Shen et al., (2021); Tian et al., (2019), and LLM-based methods Liu et al., 2024b ; Russell-Gilbert et al., (2024); Liu et al., 2024a . Prediction-based methods identify anomalies by comparing predicted values to ground truth, but they are vulnerable to historical inference errors Wang et al., (2024). Reconstruction-based methods address this limitation. However, they cannot leverage expert knowledge in specific application scenarios without model adaptation. These methods identify anomalies by learning normal patterns from training data and flagging data samples that deviate from these patterns. Yet, in scenarios where it is impossible to observe or collect all normal patterns (e.g., airplane monitoring), reconstruction-based methods are ineffective. In contrast, LLM-based methods introduce expert knowledge effectively through techniques like retrieval-augmented generation (RAG) Lewis et al., (2020) and chain of thought (COT) Wei et al., (2022) without requiring model modifications. However, as natural language-based models, LLMs are insensitive to value fluctuations and normal pattern extraction. None of these existing methods fully leverage the complementary strengths of LLM-based approaches and TSADM.
@@ A.2 Proof of Theorem 1
When $\mathcal{L}(a,b)=(a-b)^{2}$, we can transform Eq. 5 to Eq. 9, where $\lambda_{1}=\frac{D_{intra}}{D_{intra}+D_{inter}}$ and $\lambda_{2}=\frac{D_{inter}}{D_{intra}+D_{inter}}$.
@@
According to KKT condition Boyd and Vandenberghe, (2004), the optimal solution of $\mathcal{L}_{\beta}$ should satisfy Eq. 10.
@@
As shown in Appendix A.9, the gradient of architecture of conditional network can not be equal to 0. Thus, $02\lambda_{1}(y+\epsilon_{s}-\hat{S}^{*}(\mathcal{P}))+2\lambda_{2}(y+\epsilon_{S}-\hat{S}^{*}(\mathcal{P}))=$ 0. Then, $\hat{S}^{*}(\mathcal{P})=y+\lambda_{1}\epsilon_{s}+\lambda_{2}\epsilon_{S}$. Thus, $\mathbb{E}[(\hat{S}^{*}(\mathcal{P})-y)^{2}]=\mathbb{E}[(\lambda_{1}\epsilon_{s}+\lambda_{2}\epsilon_{S})^{2}]$. Since $f(x)=x^{2}$ is a convex function, according to Jensen’s Inequality, $\mathbb{E}[(\lambda_{1}\epsilon_{s}+\lambda_{2}\epsilon_{S})^{2}]\geq[\mathbb{E}(\lambda_{1}\epsilon_{s}+\lambda_{2}\epsilon_{S})]^{2}$. Since $\epsilon_{1}\sim\mathcal{D}_{s}(\mu_{s},\sigma_{s}),\epsilon_{2}\sim\mathcal{D}_{S}(\mu_{S},\sigma_{S})$, we have $[\mathbb{E}(\lambda_{1}\epsilon_{s}+\lambda_{2}\epsilon_{S})]^{2}=(\lambda_{1}\mu_{1}+\lambda_{2}\mu_{2})^{2}$. Thus, $\mathbb{E}[(\hat{S}^{*}(\mathcal{P})-y)^{2}]\geq(\lambda_{1}\mu_{1}+\lambda_{2}\mu_{2})^{2}$
@@ A.3 Proof of Lemma 1
When using stochastic gradient descent algorithm, the gradient of Eq. 5 and Eq. 6 is shown in Eq. 11 and Eq. 12 respectively, where $\mathcal{B}$ is the set of sample in batch and $|\mathcal{B}|$ is the size of set $\{(r,t)|r\in\mathcal{B},t\in\mathcal{B}\}$. Let $\mathbf{g}_{i,j}^{*}$ denote the item in Eq. 12, whose index is $(i,j)$. Let $\mathbf{g}_{i,j}$ denote the item in Eq. 11, whose index is $(i,j)$. $\mathbf{g}_{i,j}$ can be transformed to $\mathbf{g}_{i,j}^{*}+[\lambda_{1}(\epsilon_{s,i}-\epsilon_{s,j})+\lambda_{2}(\epsilon_{S,i}-\epsilon_{S,j})](\hat{S}_{i}-\hat{S}_{j})\frac{\partial(\hat{S}_{i}-\hat{S}_{j})}{\partial\mathcal{P}}$. Thus, the gradient of Eq. 5 can be deemed as the gradient of Eq. 6 add some random noise scaled by $(\hat{S}_{i}-\hat{S}_{j})\frac{\partial(\hat{S}_{i}-\hat{S}_{j})}{\partial\mathcal{P}}$ for each sample. Let $\hat{S}^{{\dagger}}$ denote the optimal solution of Eq. 6. Thus, $\mathcal{L}_{\beta}^{*}(\hat{S})\geq\mathcal{L}_{\beta}^{*}(\hat{S}^{{\dagger}}),\forall\hat{S}$, and it satisfies Assumption 5.1 in Bu et al., (2024).
@@
$\begin{split}\mathbf{g}=-\frac{1}{|\mathcal{B}|}\sum_{(i,j)\in\{(r,t)|r\in\mathcal{B},t\in\mathcal{B}\}}\lambda_{1}(\epsilon_{s,i}-\epsilon_{s,j}+y_{i}-y_{j})(\hat{S}_{i}-\hat{S}_{j})\frac{\partial(\hat{S}_{i}-\hat{S}_{j})}{\partial\mathcal{P}}\\ +\lambda_{2}(\epsilon_{S,i}-\epsilon_{S,j}+y_{i}-y_{j})(\hat{S}_{i}-\hat{S}_{j})\frac{\partial(\hat{S}_{i}-\hat{S}_{j})}{\partial\mathcal{P}}\end{split}$ (11)
$\begin{split}\mathbf{g}^{*}=-\frac{1}{|\mathcal{B}|}\sum_{(i,j)\in\{(r,t)|r\in\mathcal{B},t\in\mathcal{B}\}}\lambda_{1}(y_{i}-y_{j})(\hat{S}_{i}-\hat{S}_{j})\frac{\partial(\hat{S}_{i}-\hat{S}_{j})}{\partial\mathcal{P}}\\ +\lambda_{2}(y_{i}-y_{j})(\hat{S}_{i}-\hat{S}_{j})\frac{\partial(\hat{S}_{i}-\hat{S}_{j})}{\partial\mathcal{P}}\end{split}$ (12)
@@
Since $y\in[0,1],\hat{S}\in[0,1]$, $|y_{i}-y_{j}|\leq 1,\forall(i,j)$ and $|\hat{S}_{i}-\hat{S}_{j}|\leq 1,\forall(i,j)$. Thus, $|\mathbf{g}_{i,j}^{*}|=|(y_{i}-y_{j})(\hat{S}_{i}-\hat{S}_{j})\frac{\partial(\hat{S}_{i}-\hat{S}_{j})}{\partial\mathcal{P}}|\leq|\frac{\partial(\hat{S}_{i}-\hat{S}_{j})}{\partial\mathcal{P}}|$. Furthermore, according to Appendix. A.9, conditional network is a multi-layer perceptron with sigmoid as its activation function. Let $\mathbf{\sigma}(x)$ denote sigmoid function. Since $\frac{\partial\mathbf{\sigma}(x)}{x}=\mathbf{\sigma}(x)(1-\mathbf{\sigma}(x))$ and $10<\mathbf{\sigma}(x)<$ 1, $|\frac{\partial(\hat{S})}{\partial\mathcal{P}}|<1$. Thus, $|\frac{\partial(\hat{S}_{i}-\hat{S}_{j})}{\partial\mathcal{P}}|\leq 1$. Thus, $|\mathbf{g}_{i,j}^{*}|\leq 1$. Thus, $|\mathbf{g}^{*}|\leq\frac{1}{|\mathcal{B}|}\sum_{i,j}|\mathbf{g}_{i,j}^{*}|=1$. Since $\mathcal{L}_{\beta}^{*}$ is $L$-Lipschitz continuous, it satisfies the inequality in Eq. 13.
$\begin{split}|\mathcal{L}_{\beta}^{*}(\mathcal{P}_{1})-(\mathcal{L}_{\beta}^{*}(\mathcal{P}_{2})+{\mathbf{g}^{*}}^{T}(\mathcal{P}_{1}-\mathcal{P}_{2}))|=\\ |[\mathcal{L}_{\beta}^{*}(\mathcal{P}_{1})-\mathcal{L}_{\beta}^{*}(\mathcal{P}_{2})]+{\mathbf{g}^{*}}^{T}(\mathcal{P}_{1}-\mathcal{P}_{2})|\leq\\ L|\mathcal{P}_{1}-\mathcal{P}_{2}|+|\mathbf{g}_{i,j}^{*}|\cdot|\mathcal{P}_{1}-\mathcal{P}_{2}|\leq\\ (L+1)\cdot|\mathcal{P}_{1}-\mathcal{P}_{2}|\end{split}$ (13)
Since $L\geq 0$, $L+1>0$, our model satisfies Assumption 5.2 in Bu et al., (2024).
We can transform $\mathbb{E}(\mathbf{g}_{i,j}^{*}-\mathbf{g}_{i,j})$ step by step, as shown in Eq. 14.
$\begin{split}\mathbb{E}(\mathbf{g}_{i,j}^{*}-\mathbf{g}_{i,j})=\mathbb{E}[(\lambda_{1}(\epsilon_{s,i}-\epsilon_{s,j})+\lambda_{2}(\epsilon_{S,i}-\epsilon_{S,j}))(\hat{S}_{i}-\hat{S}_{j})\frac{\partial(\hat{S}_{i}-\hat{S}_{j})}{\partial\mathcal{P}}]\\ =\mathbb{E}[(\lambda_{1}(\epsilon_{s,i}-\epsilon_{s,j})+\lambda_{2}(\epsilon_{S,i}-\epsilon_{S,j}))](\hat{S}_{i}-\hat{S}_{j})\frac{\partial(\hat{S}_{i}-\hat{S}_{j})}{\partial\mathcal{P}}\end{split}$ (14)
Since $\mathbb{E}(\epsilon_{s,i})=\mathbb{E}(\epsilon_{s,j}),\mathbb{E}(\epsilon_{S,i})=\mathbb{E}(\epsilon_{S,j})$, we can obtain $\mathbb{E}[\lambda_{1}(\epsilon_{s,i}-\epsilon_{s,j})+\lambda_{2}(\epsilon_{S,i}-\epsilon_{S,j})]=0$. Thus, $\mathbb{E}(\mathbf{g}_{i,j}^{*}-\mathbf{g}_{i,j})=0$. $\mathbb{E}|\mathbf{g}_{i,j}^{*}-\mathbf{g}_{i,j}|^{2}$ can be calculated as shown in Eq. 15.
$\begin{split}\mathbb{E}|\mathbf{g}_{i,j}^{*}-\mathbf{g}_{i,j}|^{2}=(\hat{S}_{i}-\hat{S}_{j})^{2}(\frac{\partial(\hat{S}_{i}-\hat{S}_{j})}{\partial\mathcal{P}})^{2}\mathbb{E}[(\lambda_{1}(\epsilon_{s,i}-\epsilon_{s,j})+\lambda_{2}(\epsilon_{S,i}-\epsilon_{S,j}))^{2}]\end{split}$ (15)
As proven above, $|\hat{S}_{i}-\hat{S}_{j}|\leq 1$, thus $(\hat{S}_{i}-\hat{S}_{j})^{2}\leq 1$. Similarly $\frac{\partial(\hat{S}_{i}-\hat{S}_{j})}{\partial\mathcal{P}}\leq 1$, we can obtain $(\frac{\partial(\hat{S}_{i}-\hat{S}_{j})}{\partial\mathcal{P}})^{2}\leq 1$. Thus, we can obtain the inequality in Eq. 16.
$\begin{split}(\hat{S}_{i}-\hat{S}_{j})^{2}(\frac{\partial(\hat{S}_{i}-\hat{S}_{j})}{\partial\mathcal{P}})^{2}\mathbb{E}[(\lambda_{1}(\epsilon_{s,i}-\epsilon_{s,j})+\lambda_{2}(\epsilon_{S,i}-\epsilon_{S,j}))^{2}]\leq\\ \mathbb{E}[(\lambda_{1}(\epsilon_{s,i}-\epsilon_{s,j})+\lambda_{2}(\epsilon_{S,i}-\epsilon_{S,j}))^{2}]\end{split}$ (16)
@@
Since $\epsilon_{s,i}\sim\mathcal{D}(\mu_{s},\sigma_{s}),\epsilon_{s,j}\sim\mathcal{D}(\mu_{s},\sigma_{s})$ independently, $\mathbf{E}[(\epsilon_{s,i}-\epsilon{s,j})^{2}]=2\sigma_{s}^{2}$. Similarly, $\mathbf{E}[(\epsilon_{S,i}-\epsilon_{S,j})^{2}]=2\sigma_{S}^{2}$. After simplifying its formation, we obtain $\mathbb{E}[(\lambda_{1}(\epsilon_{s,i}-\epsilon_{s,j})+\lambda_{2}(\epsilon_{S,i}-\epsilon_{S,j}))^{2}]=2\lambda_{1}^{2}\sigma_{s}^{2}+2\lambda_{2}^{2}\sigma_{S}^{2}$. Thus, $\mathbb{E}|\mathbf{g}_{i,j}^{*}-\mathbf{g}_{i,j}|^{2}\leq 2\lambda_{1}^{2}\sigma_{s}^{2}+2\lambda_{2}^{2}\sigma_{S}^{2}$. Then, it satisfies Assumption 5.3 in Bu et al., (2024).